# Extreme value distribution for GRB prompt data

### v1

@@ Abstract
Context. Gamma-Ray Bursts (GRBs) are known to be unpredictable in time and position. A few (observationally) exceptional events have been observed, as GRB 221009A that stands out for its fluence and peak flux, being orders of magnitude higher than what measured so far.
Aims. Analyzing the observed fluence, peak flux or duration distributions typically requires one to assume some scenarios, and the consistency of the observed data with the predictions turns out to be an important model diagnostic. However, it is also of interest to model these distributions using general statistical properties that do not rely on specific model assumptions, allowing one to derive inferences only based on the consistency of the observed distributions with the hypothesis of one single population of events that generate them.
Methods. We obtained fluences, peak fluxes and durations from the catalogues of GRBs observed by the CGRO-BATSE and Fermi-GBM instruments. We selected the extreme values in slots of equal duration and modelled their distributions by the generalized extreme value (GEV) formalism. The GEV distribution is a limit distribution naturally arising when the number of observations is large and is essentially independent of the phenomena producing the observed data.
Results. The distributions of extreme values for fluences, peak fluxes and durations are consistent with being extracted from a single population of events but the fluence and peak flux recorded for GRB 221009A constitutes a striking exception. The probability to observe such an event, assuming it is a cosmological GRB, is low, with a median value of about one event per millennium for the fluence and about one event per century for the peak flux.
@@
Gamma-Ray Bursts (GRBs) have been routinely observed by high-energy instrument since their discovery by the Vela early warning satellites (Klebesadel et al., 1973). Since then, a fairly large set of spaceborne instruments have monitored the sky at times, generating catalogues with thousands of events, such as those reported by the Burst and Transient Source Experiment (CGRO-BATSE, Goldstein et al., 2013)111https://heasarc.gsfc.nasa.gov/W3Browse/cgro/bat5bgrbsp.html, Konus-Wind (e.g., Tsvetkova et al., 2017), and Fermi Gamma-Burst Monitor (Fermi-GBM, Gruber et al., 2014; von Kienlin et al., 2014; Bhat et al., 2016; von Kienlin et al., 2020; Poolakkil et al., 2021)222https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermigbrst.html.
@@
GRBs show two usually distinct emission phases, the prompt and the afterglow (Zhang, 2018). They have been classified in (at least) two different flavors, the long soft and short hard GRBs (Kouveliotou et al., 1993; Qin et al., 2013), originating from two different classes of progenitors (Zhang, 2018). However, exceptions to a classification based only on duration and hardness ratios have been singled out (see, e.g., Ahumada et al., 2021; Zhang et al., 2021; Yang et al., 2022; Dichiara et al., 2023; Levan et al., 2024; Yang et al., 2024; Yi et al., 2025).
@@
The availability of such a large collection of data has naturally allowed for a large variety of statistical analysis about, e.g., their sky distribution (Briggs et al., 1996), their duration and hardness (Kouveliotou et al., 1993), or their luminosity function (Fenimore et al., 1993; Salvaterra et al., 2012) or relations between prompt and/or afterglow parameters (Yonetoku et al., 2004; Amati, 2006; Ghirlanda et al., 2012; Dainotti et al., 2023). In this framework, the rather recent detection of GRB 221009, with its extreme properties in terms of fluence and peak flux (see, e.g., Frederiks et al., 2023; Burns et al., 2023; Lesage et al., 2023), represents a formidable test bench for any attempt to model the observed population(s) of cosmological GRBs and even the possibility of physics beyond the standard model (Galanti et al., 2024, 2025).
@@
GRB 221009A (e.g., Malesani et al., 2023, and references therein), was rapidly dubbed the “BOAT”, the “Brightest Of All Times”. Indeed, it was an exceptionally bright GRBs producing a fluence and peak flux (after correction for saturation) more than an order of magnitude greater than any other event detected so far (Burns et al., 2023, and references therein). This event is of course interesting by itself, but several studies have also been devoted to establish the probability of observing such an event if GRB 221009A were a regular cosmological GRBs. The results unavoidably depend on the various assumptions involved in the analysis. In Malesani et al. (2023) rates range from one case every few decades up to a century. Instead, Burns et al. (2023) obtained, by geometric extrapolation of the distributions for a multi-mission catalogue of long GRBs, that the fluence of GRB 221009A makes it a once-in-10,000-year event. Estimates in 1 for to years have been obtained by several more authors (e.g., Atteia et al., 2025; Lan et al., 2023; O’Connor et al., 2023; Williams et al., 2023). Such a low rate stimulated the researchers to find different possibilities to avoid the (statistical) contradiction between the actual observation and the length of the sky monitoring (a few decades). Bloom (2022) suggested that GRB 221009A could be the result of Galactic lensing. Finke & Razzaque (2024) discussed the possibility that GRB 221009A belonged to a different population of narrow-beam GRBs, so that its occurrence rate could be as high as approximately one per 200 yr. Eventually, the possibility that this event was not a cosmological GRB has also been discussed (e.g., Navia et al., 2024).
@@
The problem of determining the true rate of events comparable to GRB 221009A is therefore still open and full of consequences for our understanding of the cosmological GRB population. In this paper, we propose a different approach, almost model-independent, i.e., based only on the assumptions that the detection probability of bright GRBs by high-energy satellites is essentially constant during the lifetime of the mission and that the probability of occurrence of such an event is also constant in time, i.e., it is stationary. In fact, we modelled the set of bright events observed by the CGRO-BATSE and Fermi-GBM catalogues by the so-called generalized extreme value distribution (GEV, Beirlant et al., 2004; Coles, 2001), i.e. a statistical distribution naturally emerging asymptotically and almost independently of the phenomena under study. In Sect. 2 we describe the properties of the GEV distribution, in Sect. 3 we introduce the data set used in this work and in Sect. 4 the analysis carried out. Finally, in Section 5 we comment our results and present our conclusions.
2 The Generalized Extreme Value distribution
@@
The extreme value theory, i.e. the attempt to evaluate the probability of occurrence of extreme measurements, is a rapidly evolving field of modern statistics. It is applied to many scenarios, i.e., climatology, engineering, physics, and economy (Coles, 2001). In astronomy, recent applications include the modelling of the brightest galaxies in a survey (Heather et al., 2024), the number of sunspots (Zhang et al., 2024), the maxima of periodograms in time-series analysis (Gondhalekar et al., 2023; Süveges, 2014), density fields in cosmology (Repp & Szapudi, 2018), masses of galaxy clusters (Waizmann et al., 2012), etc.
@@
The most interesting aspect of the GEV distribution is that it can be proved that it holds asymptotically essentially independently of the probability distribution describing the phenomena that generate the observed data. The mathematical proof is somehow tedious (e.g., Beirlant et al., 2004) but not difficult, and there is some analogy with the central limit theorem that guarantees that, under rather general conditions, the limiting distribution of sample means for a large family of parental statistical distributions is a standard normal distribution (Fischer, 2011).
@@
If we collect the maxima, , of a large number of random variables, , identically distributed according to a given continuous distribution, , with , the GeV distribution can be written as:
@@
The parameter is usually called shape and it is related to the tail decay of the underlying distribution of . is the mean value of the distribution and is the scale. The shape divides the GEV family into three subfamilies. Negative shape parameters signal distributions of maxima of variables with a finite upper boundary (i.e., , while a positive or zero shape parameter indicate that there is no upper limit. The case is also known as the “Gumbel” distribution. It is defined as the limit function when , and takes the form:
@@
A GEV model is trained typically dividing a dataset into blocks of uniform length, e.g. a year, etc. Then, from each block of data, the extreme (maximum or minimum) is extracted, and finally a GEV model is fitted to the set of extreme values. Once trained, the GEV distribution can be used to calculate return levels. The return level associated with the return period , is simply the level expected to be exceeded once in all blocks of the same length . Credible intervals for any estimated quantity can be derived by the posterior distribution of the GEV parameters.
The choice of the length of the blocks used to extract extreme values is not driven by first principles, and there is a well-known bias-variance trade-off. The GEV is a limit distribution, therefore longer blocks, i.e. including a large number of elements, provide a better approximation. On the other hand, longer blocks, for any finite set of observed data, imply a reduced number of elements and unavoidably a poorer prediction with a larger variance. Smaller blocks might, on the contrary, give a worse approximation of a GEV distribution, yielding biased predictions. Therefore, the choice of block length, unless driven by physical reasons, is a trial and test process, and the results will be tested using appropriate model diagnostics as, in most cases, the quantile-quantile (QQ) plot (e.g., Coles, 2001). If the model is good, the points will closely follow a straight line with intersect 0 and slope 1 without strong systematic deviations.
We mention, finally, that a crucial assumption of extreme value analysis is that the original distribution of data where extreme values are extracted from should be stationary. If this does not happen it is indeed possible to include in the analysis exogenous variables to model the non-stationarity (e.g. trends, seasonality, etc., Coles, 2001).
@@ 3 Data set selection
In this work, we based our analyses on the Fermi-GBM GRB catalogue (e.g., Poolakkil et al., 2021) and we also considered the CGRO-BATSE GRB catalogue (Goldstein et al., 2013), as a comparison case with no events comparable to GRB 221009A. The GBM catalogue contains 4029 events from July 2008 to June 2025333Update up to June 26, 2025. , while the BATSE catalogue contains 2145 events from April 1991 to August 1996. From these catalogues, we extracted the trigger time, the fluence, the peak flux and the GRB duration, expressed, as usual, by the parameter , the duration of the time interval during which 90% of the total observed counts were detected. In more detail, from the BATSE catalogue we considered the total fluence in the energy range 20-300 keV (summing the fluences recorded in the 20-50, 50-100 and 100-300 keV bands), in units of erg cm-2 and the peak flux in the 20-300 keV energy range on the 1024 ms timescale in units of photons cm-2 sec-1. From the GBM catalogue we considered the fluence in the 10-1000 keV energy range, in erg and the peak flux in the same energy range on the 1024 ms timescale in units of photons cm-2 sec-1. Choosing different baselines for the peak flux computation does not change any of the conclusions reported in this paper.
@@
Since we are interested in the extreme values for the considered parameters, it is necessary to discuss the various biases affecting data obtained from the selected catalogues. In fact, the brightest events, are often affected by saturation. A correction for the saturation is typically carried out comparing results from different instruments by a suitable cross-calibration, and/or a modelling of the spectral evolution of the event in order to infer the expected behaviour when the considered instrument was unable to provide reliable information. In most cases, such a correction is of limited entity and not important for the present study. GRB 221009A is an exception due to its peculiar brightness. A saturation-corrected fluence was computed in Lesage et al. (2023) and the reported value is more than two times higher than the value reported in the GBM catalogue. A saturation corrected peak-flux was computed by Frederiks et al. (2023), basing on Konus-Wind data in an energy band close to the Fermi-GBM one. This estimate turns out to be an order of magnitude higher than what reported in the GBM catalogue. In the following analysis we adopted the saturation-corrected values for the BOAT.
@@
In addition, the BATSE catalogue is known to be incomplete toward the highest fluxes and fluences (see, e.g., Kaneko et al., 2006; Burns et al., 2023), due to several factors, as saturation and triggering strategy. This unavoidably affects our capability to derive a solid GEV distribution for the parameter of our interest and incompleteness in the derived extreme values can make more uncertain the extrapolation toward the highest fluences and peak fluxes, beyond the statistical uncertainties. On the other hand, as far as the bulk of the high fluence and flux reported events are reliable, the global features of the computed GEV distributions should be still of interest.
@@
Independently of the considered catalogue, events with very long duration might suffer from biases due to various possible constrains affecting instrument efficiency aboard of low orbit satellites (e.g. Earth occultation, South-Atlantic anomaly crossing, telemetry limitations, etc.). Catalogues are therefore likely truncated missing or incorrectly recording some of longest durations GRBs. The BATSE catalogue does not report events longer than s, while the GBM catalogue seems to be limited at , indicating the higher level of incompleteness suffered from the former. A few ultra-long duration GRBs () have also been singled out (e.g., Levan et al., 2014), opening a long-lasting debate about their possible different origin from the bulk of cosmological GRBs. Deriving the true duration of a very long GRB can be a difficult task, often relying on data from multiple missions. Inclusion of these data in the present analysis could undoubtedly be of interest, but would also need a proper homogenization since we should derive a duration seen by the given (BATSE or GBM) instrument without limitations, i.e., a different quantity with respect to the duration of emission activity of the source at any energy. Since the focus of the present work is on the possible peculiarity of GRB 221009A, whose duration is well within the recorded values in the GBM catalogue, we decided not to apply any correction to the duration data.

### v2

# Extreme value distribution for gamma-ray-burst prompt data

@@ Abstract
Context. Gamma-ray bursts (GRBs) are known to be unpredictable in time and position. A few (observationally) exceptional events have been observed, such as GRB 221009A, which stands out for having a fluence and peak flux orders of magnitude higher than what has been measured so far.
Aims. Analysing the observed fluence, peak flux, or duration distributions typically requires one to assume some scenarios, and the consistency of the observed data with the predictions turns out to be an important model diagnostic. However, it is also of interest to model these distributions using general statistical properties that do not rely on specific model assumptions, allowing one to derive inferences only based on the consistency of the observed distributions with the hypothesis of one single population of events that generate them.
Methods. We obtained fluences, peak fluxes, and durations from the catalogues of GRBs observed by the CGRO-BATSE and Fermi-GBM instruments. We selected the extreme values in slots of equal duration and modelled their distributions using the generalised extreme value (GEV) formalism. The GEV distribution is a limit distribution naturally arising when the number of observations is large and is essentially independent of the phenomena producing the observed data.
Results. The distributions of extreme values for fluences, peak fluxes, and durations are consistent with being extracted from a single population of events, but the fluence and peak flux recorded for GRB 221009A constitute a striking exception. The probability of observing such an event, assuming it is a cosmological GRB, is low, with a median value of about one event per millennium for the fluence and about one event per century for the peak flux.
@@
Gamma-ray bursts (GRBs) have been routinely observed by high-energy instruments since their discovery by the Vela early warning satellites (Klebesadel et al., 1973). Since then, a fairly large set of space-borne instruments have monitored the sky at times, generating catalogues with thousands of events, such as those reported by the Burst and Transient Source Experiment (CGRO-BATSE, Goldstein et al., 2013),111https://heasarc.gsfc.nasa.gov/W3Browse/cgro/bat5bgrbsp.html Konus-Wind (e.g. Tsvetkova et al., 2017), and the Fermi Gamma-Burst Monitor (Fermi-GBM, Gruber et al., 2014; von Kienlin et al., 2014; Bhat et al., 2016; von Kienlin et al., 2020; Poolakkil et al., 2021).222https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermigbrst.html
@@
Gamma-ray bursts show two usually distinct emission phases: the prompt and the afterglow (Zhang, 2018). They have been classified in (at least) two different groups, including the long-soft and the short-hard GRBs (Kouveliotou et al., 1993; Qin et al., 2013), originating from two different classes of progenitors (Zhang, 2018). However, exceptions to a classification based only on duration and hardness ratios have been singled out (see e.g. Ahumada et al., 2021; Zhang et al., 2021; Yang et al., 2022; Dichiara et al., 2023; Levan et al., 2024; Yang et al., 2024; Yi et al., 2025).
@@
The availability of such a large collection of data has naturally allowed for a large variety of statistical analysis concerning, for example, their sky distribution (Briggs et al., 1996), their duration and hardness (Kouveliotou et al., 1993), and their luminosity function (Fenimore et al., 1993; Salvaterra et al., 2012) or relations between prompt and/or afterglow parameters (Yonetoku et al., 2004; Amati, 2006; Ghirlanda et al., 2012; Dainotti et al., 2023). In this framework, the rather recent detection of GRB 221009, with its extreme properties in terms of fluence and peak flux (see e.g. Frederiks et al., 2023; Burns et al., 2023; Lesage et al., 2023), represents a formidable test bench for any attempt to model the observed population(s) of cosmological GRBs, and even the possibility of physics beyond the standard model (Galanti et al., 2024, 2025).
@@
This exceptional GRB (e.g. Malesani et al., 2023, and references therein) was rapidly dubbed the Brightest Of All Times (BOAT). Indeed, it was extremely bright GRB, producing a fluence and peak flux (after correction for saturation) more than an order of magnitude greater than any other event detected so far (Burns et al., 2023, and references therein). This event is of course interesting by itself, but several studies have also been devoted to establishing the probability of observing such an event if GRB 221009A were a regular cosmological GRB. The results unavoidably depend on the various assumptions involved in the analysis. In Malesani et al. (2023), rates range from one case every few decades up to a century. Instead, Burns et al. (2023) obtained, by geometric extrapolation of the distributions for a multi-mission catalogue of long GRBs, that the fluence of GRB 221009A makes it a once-in-ten-thousand-years event. Estimates in 1 for to years have been obtained by several other authors (e.g. Atteia et al., 2025; Lan et al., 2023; O’Connor et al., 2023; Williams et al., 2023). Such a low rate stimulated the researchers to find different ways to avoid the (statistical) contradiction between the actual observation and the length of the sky monitoring (a few decades). Bloom (2022) suggested that GRB 221009A could be the result of Galactic lensing. Finke & Razzaque (2024) discussed the possibility that GRB 221009A belonged to a different population of narrow-beam GRBs, so that its occurrence rate could be as high as approximately one per 200 yr. Eventually, the possibility that this event was not a cosmological GRB has also been discussed (e.g., Navia et al., 2024).
@@
The problem of determining the true rate of events comparable to GRB 221009A therefore remains, and it comes with many consequences for our understanding of the cosmological GRB population. In this paper, we propose a different, almost model-independent approach; i.e. based only on the assumptions that the detection probability of bright GRBs by high-energy satellites is essentially constant during the lifetime of the mission and that the probability of such an event is also constant in time (i.e. it is stationary). In fact, we modelled the set of bright events observed by the CGRO-BATSE and Fermi-GBM catalogues using the so-called generalised extreme value distribution (GEV; Beirlant et al., 2004; Coles, 2001); i.e. a statistical distribution naturally emerging asymptotically and almost independently of the phenomena under study. In Sect. 2, we describe the properties of the GEV distribution. In Sect. 3, we introduce the data set used in this work, and we discuss the analysis carried out Sect. 4. Finally, in Sect. 5, we comment our results and present our conclusions.
2 The GEV distribution
@@
The extreme value theory, i.e. the attempt to evaluate the probability of occurrence of extreme measurements, is a rapidly evolving field of modern statistics. It is applied to many scenarios such as climatology, engineering, physics, and economy (Coles, 2001). In astronomy, recent applications include the modelling of the brightest galaxies in a survey (Heather et al., 2024), the number of sunspots (Zhang et al., 2024), the maxima of periodograms in time-series analysis (Gondhalekar et al., 2023; Süveges, 2014), density fields in cosmology (Repp & Szapudi, 2018), masses of galaxy clusters (Waizmann et al., 2012), and so on.
@@
The most interesting aspect of the GEV distribution is that it can be proved that it holds asymptotically, essentially independently of the probability distribution describing the phenomena that generate the observed data. Obtaining mathematical proof is somewhat tedious (e.g. Beirlant et al., 2004), but not difficult. There is some analogy with the validity of the central limit theorem, i.e. that under rather general conditions the limiting distribution of sample means for a large family of parental statistical distributions is a standard normal distribution (Fischer, 2011).
@@
If we collect the maxima, , of a large number of random variables, , identically distributed according to a given continuous distribution, , with , the GeV distribution can be written as
@@
The parameter is usually called shape, and it is related to the tail decay of the underlying distribution of . is the mean value of the distribution and is the scale. The shape divides the GEV family into three sub-families. Negative shape parameters signal distributions of maxima of variables with a finite upper boundary (i.e., , while a positive or zero shape parameter indicates that there is no upper limit. The case is also known as the ‘Gumbel’ distribution. It is defined as the limit function when , and it takes the form
@@
A GEV model is typically trained by dividing a dataset into blocks of uniform length (e.g. a year, etc.). Then, from each block of data, the extreme (maximum or minimum) is extracted, and finally a GEV model is fitted to the set of extreme values. Once trained, the GEV distribution can be used to calculate return levels. The return level associated with the return period, , is simply the level expected to be exceeded once in all blocks of the same length . Credible intervals for any estimated quantity can be derived by the posterior distribution of the GEV parameters.
The choice of the length of the blocks used to extract extreme values is not driven by first principles, and there is a well-known bias-variance trade-off. The GEV is a limit distribution; therefore, longer blocks –i.e. including a large number of elements– provide a better approximation. On the other hand, longer blocks, for any finite set of observed data, imply a reduced number of elements and unavoidably a poorer prediction with a larger variance. Smaller blocks might, on the contrary, give a worse approximation of a GEV distribution, yielding biased predictions. Therefore, the choice of block length, unless driven by physical reasons, requires a trial-and-improvement process; the results are tested using appropriate model diagnostics such as, in most cases, the quantile-quantile (QQ) plot (e.g. Coles, 2001). If the model is good, the points will closely follow a straight line with intersect 0 and slope 1 without strong systematic deviations.
We mention, finally, that a crucial assumption of extreme value analysis is that the original distribution of data that extreme values are extracted from should be stationary. If this does not happen, it is indeed possible to include exogenous variables in the analysis to model the non-stationarity values (e.g. trends, seasonality, etc.; Coles, 2001).
@@ 3 Data set selection
In this work, we based our analyses on the Fermi-GBM GRB catalogue (e.g. Poolakkil et al., 2021), and we considered the CGRO-BATSE GRB catalogue (Goldstein et al., 2013) as a comparison case with no events comparable to GRB 221009A. The GBM catalogue contains 4029 events from July 2008 to June 2025,333Updated up to June 26 2025. while the BATSE catalogue contains 2145 events from April 1991 to August 1996. From these catalogues, we extracted the trigger time, the fluence, the peak flux, and the GRB duration; the latter is expressed, as usual, by the parameter , the duration of the time interval during which 90% of the total observed counts were detected. In more detail, from the BATSE catalogue we considered the total fluence in the energy range 20-300 keV (summing the fluences recorded in the 20-50, 50-100, and 100-300 keV bands), in units of erg cm-2 and the peak flux in the 20-300 keV energy range on the 1024 ms timescale in units of photons cm-2 sec-1. From the GBM catalogue we considered the fluence in the 10-1000 keV energy range, in erg cm-2 and the peak flux in the same energy range on the 1024 ms timescale in units of photons cm-2 sec-1. Choosing different baselines for the peak flux computation does not change any of the conclusions reported in this paper.
@@
Since we are interested in the extreme values for the considered parameters, it is necessary to discuss the various biases affecting data obtained from the selected catalogues. In fact, the brightest events, are often affected by saturation. A correction for the saturation is typically carried out comparing results from different instruments by a suitable cross-calibration, and/or a modelling of the spectral evolution of the event in order to infer the expected behaviour when the considered instrument was unable to provide reliable information. In most cases, such a correction is of limited entity and not important for the present study. GRB 221009A is an exception due to its peculiar brightness. A saturation-corrected fluence was computed in Lesage et al. (2023), and the reported value is more than two times higher than the value reported in the GBM catalogue. A saturation-corrected peak flux was computed by Frederiks et al. (2023) based on Konus-Wind data in an energy band close to the Fermi-GBM one. This estimate turns out to be an order of magnitude higher than what was reported in the GBM catalogue. In the following analysis we adopted the saturation-corrected values for the BOAT.
@@
In addition, the BATSE catalogue is known to be incomplete towards the highest fluxes and fluences (see e.g. Kaneko et al., 2006; Burns et al., 2023) due to several factors including saturation and triggering strategy. This unavoidably affects our capability to derive a solid GEV distribution for the parameter of our interest, and incompleteness in the derived extreme values can make the extrapolation towards the highest fluences and peak fluxes more uncertain beyond the statistical uncertainties. On the other hand, as far as the bulk of the high-fluence and -flux reported events are reliable, the global features of the computed GEV distributions should be still of interest.
@@
Independently of the considered catalogue, events with very long duration might suffer from biases due to various possible constrains affecting instrument efficiency aboard of low orbit satellites (e.g. Earth occultation, South-Atlantic anomaly crossing, telemetry limitations, etc.). Catalogues are therefore likely truncated, missing, or incorrectly record some of the longest duration GRBs. The BATSE catalogue does not report events longer than s, while the GBM catalogue seems to be limited at , indicating the higher level of incompleteness suffered by the former. A few ultra-long-duration GRBs () have also been singled out (e.g. Levan et al., 2014), opening a long debate about their possible different origin compared to the bulk of cosmological GRBs. Deriving the true duration of a very long GRB can be a difficult task, often relying on data from multiple missions. Inclusion of these data in the present analysis could undoubtedly be of interest, but it would also require proper homogenisation since we should derive a duration seen by the given (BATSE or GBM) instrument without limitations; i.e. a different quantity with respect to the duration of emission activity of the source at any energy. Since the focus of the present work is on the possible peculiarity of GRB 221009A, whose duration is well within the recorded values in the GBM catalogue, we decided not to apply any correction to the duration data.